Draw the shear and moment diagrams. 4 kN/m А E B ip 2 m 3 m 1 m 2 m 15 kN 10 kN
Answer #1

IkN/m to im 2m 3m 2m 10 15 -4x2= -8 - 4X2 +15=7 Vc shear VA=0 VB (left) VB (Right) - -4X5+15 2-5 VD = -4x5 +15=-5 Creff) XD -pls do appreciate it by giving it a thumbs up. Thanks